### v5

# Analytic Scalar Field Theory

Abstract
We discuss scalar field theories with potentials $V\left(\phi\right)=\kappa\left(\phi^{2}\right)^{\nu}$ for generic $\nu$. We conjecture that these models evade various no-go theorems for scalar field theories in four spacetime dimensions.
In four-dimensional spacetime (4D) the two simplest renormalizable scalar field theories are well-known to be flawed. Consider the standard Lagrangian density
$L=\frac{1}{2}\left(\partial\phi\right)^{2}-\frac{1}{2}m^{2}\phi^{2}-V\left(\phi\right)$ (1)
If $V\left(\phi\right)=\kappa\phi^{3}$, the theory has no ground state, i.e. the energy density is unbounded below, as is evident from consideration of semi-classical coherent states [1]. If $V\left(\phi\right)=\kappa\phi^{4}$ for real coupling constant $\kappa>0$, the theory has a ground state, by the same considerations, but it is “trivial” due to renormalization effects in 4D, i.e. only $\kappa=0$ is sensible under renormalization [2, 3, 4].
Moreover, if $V$ involves higher integer powers of $\phi$ in 4D the theory is non-renormalizable by well-known power counting arguments. As a consequence of all this the general folklore is: In 4D there are no consistent and tractable interacting quantum field theories involving only a scalar field.
Nevertheless, almost 60 years ago there was an intriguing suggestion by Delbourgo, Salam, and Strathdee [5] that non-polynomial interactions could perhaps lead to interesting and non-trivial renormalizable quantum field theories involving only scalar fields, even in 4D, albeit at considerable computational expense. More recently the general folklore has also been called into question by proponents of PT-symmetric theories [6], but again at a high cost in terms of computational methods.
As argued here, the aforementioned issues and their possible work-arounds can be considered in a tractable framework through the use of specific analytic functions of $\phi$, with computations more manageable than encountered in previous work. The models proposed here are less challenging than the generic situations studied in [5], while PT-symmetric systems are essentially a special case after only minor modifications.
[…]
To deal with an arbitrary power of $\phi^{2}$ the basic idea is to define $V\left(\phi\right)=\kappa\left(\phi^{2}\right)^{\nu}$ as a linear superposition of Gaussians. That is to say, use the exact mathematical result
$\left(\phi^{2}\right)^{\nu}=\Gamma\left(1+\nu\right)\frac{i}{2\pi}{\displaystyle\oint}\left(-z\right)^{-1-\nu}\exp\left(-z\phi^{2}\right)dz$ (2)
where $\phi^{2}$ is assumed to be real and positive, and where the integration $\oint$} is along a Hankel contour around the positive $z$ half-line [7, 8]. The right-hand side of (2) is well-defined and completely unambiguous so long as $\nu$ is not a negative integer. In particular, a “marginal” interaction in $N$ spacetime dimensions (ND) is given by $\nu=N/\left(N-2\right)$. In ND the coupling constant $\kappa$ is always dimensionless for potential energy density $V\left(\phi\right)=\kappa\left(\phi^{2}\right)^{\frac{N}{N-2}}$.
As an illustration of the method, consider the $O\left(\kappa\right)$ contribution to the vacuum expectation value (VEV) of the potential energy density. To this order in $\kappa$, only the free-field VEV of a Gaussian is required to complete the calculation. As shown in the Appendix this is
$\left\langle\exp\left(-z\phi^{2}\right)\right\rangle=\frac{1}{\sqrt{1+2z\left\langle\phi^{2}\right\rangle}}$ (3)
The calculation therefore reduces to the free-field evaluation of $\left\langle\phi^{2}\right\rangle$, which can be done in ND using a momentum cut-off, or $\zeta$ function regularization, or dimensional regularization. All that remains is evaluation of the manifestly convergent contour integral in (2). Thus, by rescaling the variable of integration,
$\left\langle\left(\phi^{2}\right)^{\nu}\right\rangle=\Gamma\left(1+\nu\right)\frac{i}{2\pi}{\displaystyle\oint}\left(-z\right)^{-1-\nu}\frac{1}{\sqrt{1+2z\left\langle\phi^{2}\right\rangle}}dz=\left\langle\phi^{2}\right\rangle^{\nu}\Gamma\left(1+\nu\right)\frac{i}{2\pi}{\displaystyle\oint}\left(-z\right)^{-1-\nu}\frac{1}{\sqrt{1+2z}}dz$ (4)
where $\left\langle\phi^{2}\right\rangle$ is assumed to be positive. The integral yields to routine contour evaluation methods that take into account the analytic structure of the integrand. The result is
$\frac{i}{2\pi}{\displaystyle\oint}\left(-z\right)^{-1-\nu}\frac{1}{\sqrt{1+2z}}dz=\frac{2^{\nu}}{\sqrt{\pi}}\frac{\Gamma\left(\nu+1/2\right)}{\Gamma\left(\nu+1\right)}$ (5)
The final result for the VEV in question is then
$\left\langle V\left(\phi\right)\right\rangle=\kappa\left\langle\left(\phi^{2}\right)^{\nu}\right\rangle=\frac{2^{\nu}}{\sqrt{\pi}}~\Gamma\left(\nu+1/2\right)~\kappa\left\langle\phi^{2}\right\rangle^{\nu}$ (6)
The coefficient in this result is just an analytic continuation of the double factorial for odd integers, $\left(2\nu-1\right)!!$.
So, at least to this order in $\kappa$, quantum effects have only modified the form expected from classical field theory by a simple replacement of $\kappa$ with
$\kappa_{\operatorname*{effective}}=2^{\nu}\Gamma\left(\nu+1/2\right)~\kappa/\sqrt{\pi}$ (7)
where all the $\nu$ dependence is exact and explicit. Note that (6) is valid even for some negative powers of $\phi^{2}$, at least so long as $\nu>-1/2$ if the coefficient of the VEV is to be finite and positive.
Similar results follow for the $O\left(\kappa\right)$ corrections to the scalar propagator, which are again exact as functions of $\nu$, as can be obtained just by differentiation of $\left\langle V\left(\phi\right)\right\rangle$. To this order the effect is only a shift in the mass of the field as is evident upon summing the usual geometric series involving the $O\left(\kappa\right)$ correction.
$\Delta\left(k\right)=\frac{i}{k^{2}-m^{2}-2\nu\kappa_{\operatorname*{effective}}\left\langle\phi^{2}\right\rangle^{\nu-1}}$ (8)
Momentum dependent corrections to $\Delta\left(k\right)$ begin at order $\kappa^{2}$. These will be discussed elsewhere [9]. However, the method to compute higher order corrections is a straightforward procedure based on (2).
For example, the order $\kappa^{2}$ computation of $\left\langle V\left(\phi\right)\right\rangle$ follows from $\left\langle\left(\phi^{2}\left(0\right)\right)^{\nu}\left(\phi^{2}\left(x\right)\right)^{\nu}\right\rangle$ upon replacing the powers of $\phi^{2}$ with superpositions of Gaussians as defined in (2). Evaluation of the VEV is then facilitated by the identity
$\left\langle e^{-w\phi^{2}\left(0\right)}e^{-z\phi^{2}\left(x\right)}\right\rangle=\frac{1}{\sqrt{1+\left(2w+2z\right)\left\langle\phi^{2}\left(0\right)\right\rangle+4wz\left(\left\langle\phi^{2}\left(0\right)\right\rangle^{2}-\left\langle\phi\left(x\right)\phi\left(0\right)\right\rangle^{2}\right)}}$ (9)
This identity, as well as that for higher multinomials involving Gaussians, follows from a Stratonovich-Hubbard transformation [10, 11], or else from standard Wick contraction combinatorics [12], as given by
$\left\langle e^{-w\phi^{2}\left(0\right)}\phi^{2n}\left(x\right)\right\rangle=\frac{\left\langle\phi^{2}\left(0\right)\right\rangle^{n}}{\left(1+2w\left\langle\phi^{2}\left(0\right)\right\rangle\right)^{1/2}}\sum_{k=0}^{n}\dbinom{2n}{2k}\frac{\left(2n-2k-1\right)!!\left(2k-1\right)!!}{\left(1+2w\left\langle\phi^{2}\left(0\right)\right\rangle\right)^{k}}\left(-2w\frac{\left\langle\phi\left(x\right)\phi\left(0\right)\right\rangle^{2}}{\left\langle\phi^{2}\left(0\right)\right\rangle}\right)^{k}$ (10)
The generating functional for the latter VEVs is then obtained by taking $\sum_{n=0}^{\infty}\frac{\left(-z\right)^{n}}{n!}\phi^{2n}\left(x\right)$ to obtain (9).
For this example the integration over Hankel contours once again yields to routine evaluation methods. The final result is a Gauss hypergeometric function.
$\kappa^{2}\left\langle\left(\phi^{2}\left(0\right)\right)^{\nu}\left(\phi^{2}\left(x\right)\right)^{\nu}\right\rangle=\kappa_{\text{effective}}^{2}\left\langle\phi^{2}\left(0\right)\right\rangle^{2\nu}\left.{}_{2}F_{1}\left(-\nu,-\nu;~\frac{1}{2};~\frac{\left\langle\phi\left(x\right)\phi\left(0\right)\right\rangle^{2}}{\left\langle\phi^{2}\left(0\right)\right\rangle^{2}}\right)\right$. (11)
where again $\kappa_{\text{effective}}$ is given by (7). It is satisfying that this result behaves in accord with the cluster decomposition property [13] in the sense that $\kappa^{2}\left\langle\left(\phi^{2}\left(0\right)\right)^{\nu}\left(\phi^{2}\left(x\right)\right)^{\nu}\right\rangle\underset{\left\langle\phi\left(x\right)\phi\left(0\right)\right\rangle\rightarrow 0}{\longrightarrow}\left(\kappa\left\langle\left(\phi^{2}\left(0\right)\right)^{\nu}\right\rangle\right)^{2}$.
A special case of interest is given by
$\kappa^{2}\left\langle\left(\phi^{2}\left(0\right)\right)^{3/2}\left(\phi^{2}\left(x\right)\right)^{3/2}\right\rangle=\frac{8}{\pi}\kappa^{2}\left\langle\phi^{2}\left(0\right)\right\rangle^{3}\left.{}_{2}F_{1}\left(-\frac{3}{2},-\frac{3}{2};~\frac{1}{2};~\frac{\left\langle\phi\left(x\right)\phi\left(0\right)\right\rangle^{2}}{\left\langle\phi^{2}\left(0\right)\right\rangle^{2}}\right)\right$. (12)
where this particular $\left.{}_{2}F_{1}\right$. reduces to elementary functions.
$\left.{}_{2}F_{1}\left(-\frac{3}{2},-\frac{3}{2};~\frac{1}{2};~z^{2}\right)\right$. $=\frac{1}{4}\left(\left(4+11z^{2}\right)\sqrt{1-z^{2}}+3\left(3+2z^{2}\right)z\arcsin z\right)$
$=1+\frac{9}{2}z^{2}+\frac{3}{8}z^{4}+\frac{1}{80}z^{6}+\frac{9}{4480}z^{8}+O\left(z^{10}\right)$ (13)
Compare (12) to that of conventional $\kappa\phi^{3}$ theory [1] for which there are only two terms in the corresponding VEV, rather than an infinite series, namely,
$\kappa^{2}\left\langle\phi\left(0\right)^{3}\phi\left(x\right)^{3}\right\rangle=\kappa^{2}\left(9\left\langle\phi^{2}\left(0\right)\right\rangle^{2}\left\langle\phi\left(x\right)\phi\left(0\right)\right\rangle+6\left\langle\phi\left(x\right)\phi\left(0\right)\right\rangle^{3}\right)$ (14)
The comparison highlights the difference between $V=\kappa\phi^{3}$ and $V=\kappa\left(\phi^{2}\right)^{3/2}$ as defined by (2).
The above results raise the possibility of scalar field theories with non-zero $\left\langle\phi^{2}\right\rangle$ through more exotic mechanisms than usually encountered in the literature. Consider a potential defined as
$V\left(\phi\right)=\frac{\kappa_{1}}{\left(\phi^{2}\right)^{\nu_{1}}}+\kappa_{2}\left(\phi^{2}\right)^{\nu_{2}}$ (15)
with real and positive $\kappa_{1}$, $\kappa_{2}$, $\nu_{1}$, & $\nu_{2}$. With the additional assumption $0<\nu_{1}<1/2$, free-field VEVs of both terms in $V$ give an expectation value to first order in either $\kappa$ as
$\left\langle V\left(\phi\right)\right\rangle=\frac{\kappa_{1\text{ }\operatorname*{effective}}}{\left\langle\phi^{2}\right\rangle^{\nu_{1}}}+\kappa_{2\text{ }\operatorname*{effective}}\left\langle\phi^{2}\right\rangle^{\nu_{2}}$ (16)
where the effective $\kappa_{1}$ and $\kappa_{2}$ are given by
$\kappa_{1\text{ }\operatorname*{effective}}=\frac{2^{\nu_{1}}\Gamma\left(1-2\nu_{1}\right)}{\Gamma\left(1-\nu_{1}\right)}~\kappa_{1}\ ,\ \ \ \kappa_{2\text{ }\operatorname*{effective}}=\frac{\Gamma\left(1+2\nu_{2}\right)}{2^{\nu_{2}}\Gamma\left(1+\nu_{2}\right)}~\kappa_{2}$ (17)
[…]
$V\left(\phi\right)$, in green, & $\left\langle V\right\rangle$ versus $\sqrt{\left\langle\phi^{2}\right\rangle}$, in orange, with minima as indicated.
Several parallels to the earlier work of Salam et al. [5] would result from power series expansions in $\phi$ about either of the minima shown.
The interacting models, as defined here, are in one respect similar to that for a 2D sine-Gordon model [14] with $V\left(\phi\right)=\kappa\cos\left(\phi\right)$, where $\kappa$ has mass dimension two. That is to say, perturbative results to any finite order in $\kappa$ [15] involve closed $n$-loop Feynman diagrams for all integer $n\geq 1$. However, in contrast to the 2D sine-Gordon model for which $\phi$ has mass dimension $\left[\phi\right]=0$, if the models considered here are to be in $N$ spacetime dimensions, then $\left[\phi\right]=\left(N-2\right)/2$. Hence the coupling constant for $V\left(\phi\right)=\kappa\left(\phi^{2}\right)^{\nu}$ has mass dimension $\left[\kappa\right]=2\nu+\left(1-\nu\right)N$. Conventional power-counting arguments, taken naively at face value, would suggest the model is super-renormalizable for $\left[\kappa\right]>0$, renormalizable for $\left[\kappa\right]=0$, and non-renormalizable for $\left[\kappa\right]<0$. To see whether these naive suggestions are correct is one objective of future studies [9].
Appendix: Some combinatorics
The free field VEV of a Gaussian follows from
$\left\langle\exp\left(-z\phi^{2}\right)\right\rangle=\sum_{n=0}^{\infty}\frac{\left(-z\right)^{n}}{n!}\left\langle\phi^{2n}\right\rangle$ (A1)
where $\left\langle\phi^{2n}\right\rangle$ is a “flower” graph with $n$ “petals” as sketched below, each petal representing the VEV $\left\langle\phi^{2}\right\rangle$.
A Feynman flower diagram.
The combinatoric factor for this graph is $\left(2n-1\right)!!=\left(2n\right)!/\left(2^{n}n!\right)$ where $\left(2n\right)!$ corresponds to the number of ways to select each individual $\phi$ from $\phi^{2n}$, each petal has a symmetry factor of $1/2$, and the additional $1/n!$ is a symmetry factor that corresponds to the number of permutations of the petals. That is to say,
$\left\langle\phi^{2n}\right\rangle=\left(2n-1\right)!!\left\langle\phi^{2}\right\rangle^{n}$ (A2)
This immediately gives the result (3) in the text, namely,
$\left\langle\exp\left(-z\phi^{2}\right)\right\rangle=\sum_{n=0}^{\infty}\frac{\left(2n-1\right)!!}{n!}\left(-z\left\langle\phi^{2}\right\rangle\right)^{n}=\frac{1}{\sqrt{1+2z\left\langle\phi^{2}\right\rangle}}$ (A3)
This result also follows from standard operator methods, e.g., as often used to prove the BCH formula.
As explained in the text following (9), similar combinatorics, as well as other methods, give the result for the VEV of two or more Gaussians.